OpenStudy (anonymous):

Any ideas about this one? The answers I got have not worked: Henry and Irene working together can wash all the windows of their house in 5 h 52 min. Working alone, it takes Henry 4 1 2 h more than Irene to do the job. How long does it take each person working alone to wash all the windows?

OpenStudy (anonymous):

it takes him 4 1/2 more hours

OpenStudy (anonymous):

you are familiar with the formula \[\large \frac{t}{x} + \frac{t}{y} = 1\] where t = total time working together

OpenStudy (anonymous):

I've done this problem but I cannot get the correct answer. I've been on this problem for a long time

OpenStudy (anonymous):

convert 5 h 52 mins and 4 1/2 hours into minutes first what do you have?

OpenStudy (anonymous):

352/60 and 270/60

OpenStudy (anonymous):

why is there over 60? no over 60..

OpenStudy (anonymous):

so total time = 352 minutes let x = Irene's time Henry = 270 + x can you do the equation now?

OpenStudy (anonymous):

henry is 82 mins

OpenStudy (anonymous):

uhh no...you cannot just simply subtract 270 from 352..do the equation

OpenStudy (anonymous):

\[\large \frac{t}{x} + \frac{t}{y} = 1\]